An ion source is producing 6Li ions, which have mass 9.99 ✕ 10−27 kg and charge +e. The ions are accelerated by a potential difference of 15 kV and pass horizontally into a region in which there is a uniform vertical magnetic field of magnitude B = 2.4 T. Calculate the strength of the smallest electric field, to be set up over the same region, that will allow the 6Li ions to pass through undeflected.
Answer #1

change in kinetic energy = change in electrical potential energy = (potential difference)(charge)
(1/2)mv² = Vq
(1/2)(9.99e-27 kg)(v²) = (15000 V)(1.60e-19 C)
v = 6.931e5 m/s

net force = force on moving charges due to magnetic field
ma = qvB (v and B are perpendicular to each other)
m(v²/r) = qvB (centripetal acceleration)
r = mv/(qB)

r = (9.99e-27 kg)(6.931e5 m/s)/[(1.60e-19 C)(2.40 T)]
r =1.8e-2 m

d = 2r = 3.6e-2 m